OpenStudy (anonymous):

Solve for x: 16x2 - 81 = 0

OpenStudy (tkhunny):

Can you say "Difference of Squares"? Do it!

OpenStudy (idealist10):

16x^2-81=0 (4x-9)(4x+9)=0 4x-9=0 4x+9=0 4x=9 4x=-9 x=9/4, -9/4.

OpenStudy (idealist10):

You need to factor out.